OpenStudy (anonymous):

evaluate the following indefinite integral 15x^3/2dx

OpenStudy (inkyvoyd):

\(\Huge \int\frac{15x^3}{2}dx=\frac{15}{2}\int x^3 dx\) does that help any?

OpenStudy (anonymous):

its actually 15x and the exponent is 3/2

OpenStudy (inkyvoyd):

Oh.

OpenStudy (anonymous):



OpenStudy (goformit100):

First of all for solving this question do you know the conception of : power function and its formula of its integral

OpenStudy (anonymous):

sorry that's awful haha

OpenStudy (inkyvoyd):

\(\Huge \int 15x^{\frac{3}{2}}dx\)

OpenStudy (anonymous):



OpenStudy (inkyvoyd):

\(\Huge \int x^{n}dx=\frac{x^{n+1}}{n+1}+C\) for all n except n=-1

OpenStudy (inkyvoyd):

For this case, n=3/2, so just use this rule (the power rule) to obtain the answer. You also have to multiply the answer by 15 because of that 15 in front of the x.

OpenStudy (anonymous):

its 15x^ / 2d

OpenStudy (anonymous):

15x^2 / 2d

OpenStudy (inkyvoyd):

@Math_Hater- the dx is part of the notation https://en.wikipedia.org/wiki/Integral